# Unifying Collisional Models and the Monte Carlo Metropolis Method: Algorithms for Dynamics of Open Quantum Systems

### v1

@@ Abstract
Classical systems placed in contact with a thermal bath will inevitably equilibrate to a thermal state at the bath temperature. The same is not generally true for open quantum systems, which place additional conditions on the structure of the bath and system-bath interaction if thermalization is to occur. Collisional models, or repeated interaction schemes, are a category of microscopic open quantum system models that have seen growing use in studying quantum thermalization, in which the bath is modeled as a large ensemble of identical ancilla systems that sequentially interact with the system. We demonstrate that, when each bath ancilla is prepared in a thermal state with a discrete spectrum that matches the energy eigenstate transitions of the system, the system dynamics generated by the collisional model framework are identical to those generated under the Metropolis algorithm. This equivalence holds not just in the steady state regime, but also in the transient regime. As the Metropolis scheme does not require explicitly modeling the system-bath interaction, this allows it to be used as a computationally efficient alternative for simulating collisional model dynamics.
@@ I Introduction
The mechanisms and conditions under which a many-body quantum system will thermalize is a question of significant interest that bridges the fields of quantum thermodynamics, condensed matter physics, atomic, molecular, and optical physics, and quantum information. In the context of isolated (closed) quantum systems thermalization of a non-integral many-body quantum system can be achieved under the conditions of the eigenstate thermalization hypothesis [1, 2, 3, 4]. However, these conditions are far from guaranteed, and there exists a class of systems that resist thermalization, a phenomena referred to as many-body localization [5, 6].
In the context of open quantum systems it is well established that a quantum system coupled to a heat bath environment evolving under the Markovian Lindblad master equation will equilibrate to a thermal state at the temperature of the bath [7]. While the Lindblad equation is often tractable, it relies on strong assumptions about the dynamics of both the system and environment as well as the system-environment coupling, namely the Born-Markov and rotating wave approximations [7].
In recent years another approach to modeling open quantum systems has seen growing use, especially in the field of quantum thermodynamics, known as collisional models, or repeated interaction schemes [8]. In the collisional model approach the environment is assumed to consist of a collection of many identical subsystems, referred to as ancillae. The interaction between the system and environment occurs as a series of discrete unitary interactions (“collisions”) between the system and one ancilla of the environment. After the interaction, the ancilla is discarded (traced out) and a fresh ancilla is introduced at the next time step.
By microscopically modeling the system-environment interaction, collisional models have proven particularly useful in studying non-Markovian dynamics (by allowing for correlated ancillae or ancilla-ancilla interactions) [9, 10, 11, 12, 13, 14] and non-equilibrium dynamics [15, 16, 17, 18]. Notably, the typical Landblad master equation can be derived from the collisional model framework under the assumptions of non-interacting, uncorrelated ancillae [19, 20, 21, 8].
@@
Distinct from the deterministic evolution in the collisional models, the Monte Carlo methods are another class of methods used to study the thermalization of quantum systems [28, 29]. These methods rely upon stochastic sampling of states in the Hilbert space such that the dominant contributions to the ground or the thermal states are captured. These are iterative algorithms wherein each iteration, a random change to the state in the present iteration is considered, thereby generating a trial state. This trial state is either accepted or rejected in the next iteration based on the relative probability of the two states. The trial state is proposed based on some update to the present state only, making this a Markovian process, also referred to as a Markov chain. By stochastic averaging over different chains, the underlying probability distribution can be built. The Markov chain must satisfy the principles of ergodicity and detailed balance to converge to the desired distribution. Here ergodicity implies that every chosen state can be connected to any other state through a finite number of Markov chain moves, to ensure a complete sampling of states. Such Monte Carlo techniques are utilized in understanding the thermodynamic properties of lattice systems [30, 31] and can also be interpreted as a dynamic process related to the Glauber Kinetic Ising models [32, 33].
Studying open system dynamics has significant practical importance, as very few systems are truly isolated from their environment. Collisional models are conceptually important as they provide a microscopic framework that can operate outside of common assumptions such as weak system-bath interactions. This makes collisional models useful in wide range of contexts including for studying open dynamics in strongly-correlated models [15, 34], quantum optics and simulating light-matter interactions [20, 35], and modeling noise in quantum devices [14]. A notable drawback of the collisional model approach is the need to operate in the joint system-bath Hilbert space, which can become computationally unwieldy, especially for large-dimensional bath ancillae. Here we perform a detailed comparison to show that, subject to certain conditions, when modeling open system thermalization this problem can be circumvented by using a Metropolis Monte Carlo approach that exactly replicates the collisional model dynamics, in both the transient and steady state regimes, without needing to account for the Hilbert space of the bath.
@@ II Collisional model thermalization
This process, free evolution followed by interaction with a fresh environment ancilla, is then repeated for $n$ time steps. In Fig. 1 we provide a conceptual illustration of the collisional model approach.
@@ II.1 Collisional model thermalization of the XXZ model
$H_{XXZ}=-J\sum_{j=1}^{N-1}\left(\sigma^{x}_{j}\sigma^{x}_{j+1}+\sigma^{y}_{j}\sigma^{y}_{j+1}+\Delta\sigma^{z}_{j}\sigma^{z}_{j+1}\right)+\frac{h}{2}\sum_{j=1}^{N}\sigma^{z}_{j}$ (4)
@@
$\begin{split}&E_{1}=-h-J\Delta,\quad\,E_{2}=J(\Delta-2),\\ &E_{3}=h-J\Delta,\qquad E_{4}=J(2+\Delta).\end{split}$ (5)
@@
Combining Eqs. (8), (9), and (6) in Eq. (1), we numerically simulate a 20 time step collisional model for the two-site XXZ chain. We assume the initial system density matrix to be the maximally mixed state. In Fig. 3 we plot the occupation probabilities of each energy eigenstate as a function of the time step, $n$. We see that the occupation probabilities rapidly approach the thermal state values, indicating thermalization.
@@ III Thermalization under the Monte Carlo Metropolis algorithm
Next, we construct a Monte Carlo algorithm with Metropolis [36] updating to demonstrate thermalization of the XXZ model in the same context as the collisional model. Our approach is closely related to the Monte-Carlo wave function method [37, 38, 39] and the quantum trajectories method [40] used in quantum optics. There are also generalizations of the Metropolis scheme to the quantum regime that exploit circuit-based designs [41].
@@
The Metropolis updating technique [36] was introduced as an algorithm to significantly improve Monte Carlo convergence. The goal of the Metropolis scheme is to generate a sequence of states such that the distribution of these states closely resembles the desired distribution. The key feature of this sampling procedure is using the Metropolis filter function when deciding to accept or reject a proposed move. There are two steps involved in this method. The first one is to propose a move from the present state $x$ to $x^{\prime}$, which is based on the conditional proposal probability $G(x^{\prime}|x)$. The next step involves the acceptance probability $A(x^{\prime}|x)$ which then determines the acceptance of the proposed move. The detailed balance condition requires, $A(x^{\prime}|x)G(x^{\prime}|x)P(x)=A(x|x^{\prime})G(x|x^{\prime})P(x^{\prime})$, where $P(x)$ is the underlying probability distribution to be sampled. Note that the normalization factor in $P(x)$ gets canceled and thus the stochastic averaging can be done without explicitly computing the normalization factor, one key advantage of the Metropolis method. In the context of thermalization, $P(x)$ is the Gibbs distribution, with the partition function as the normalization factor. Thus the accept and reject criteria are based on the energy difference between the proposed and the current state. For classical systems, for instance, spin systems, the updates are often local, involving the flipping of the spin at the chosen site. However, in general, the update scheme can also be non-local [42, 43]. Since the spin basis states are the eigenstates of the classical spin Hamiltonians, these “classical” updates are sufficient to produce thermalization.
@@
In the context of quantum systems, the eigenvectors are, in general, non-trivial superpositions of the spin basis states. In this case, the classical spin updates are not sufficient to achieve thermalization. To do so, the update scheme must ensure the algorithm explores all the system eigenstates. This can be done using eigenstate jump operators, whose action produces a jump from the present eigenstate to any other eigenstate. The simplest choice for the proposal probability $G(x^{\prime}|x)$ is a uniform distribution, meaning that all possible “jumps” from the present eigenstate to other eigenstates will be proposed with equal probability. The Markov chain constructed in this way is a sequence of eigenstates weighted by the thermal distribution. Consequently, the low energy states are more probable in the Markov chain in comparison to the high energy states.
@@
For studying thermalization, we consider an update scheme based on the jumps between the different eigenstates. The steps of the algorithm are outlined in Algorithm 1. The first step involves computing the eigenvectors $\{\psi_{i}\}$ and eigenvalues $\{E_{i}\}$ of the XXZ Hamiltonian given in Eq. 4 for a system of size $N$. The set of eigenvectors and eigenvalues can then be used for the Metropolis updating scheme. We consider one eigenstate, $\psi_{i}$, chosen randomly from an uniform distribution, and then consider a jump to some other eigenstate $\psi_{k}$, also chosen randomly. We then use the Metropolis condition to decide whether the jump from $\psi_{i}$ to $\psi_{k}$ is accepted. These steps are then repeated $n$ times, and the corresponding time series of states constitutes one sample, also referred to as a Monte Carlo run. The time-dependent density matrix is then constructed by averaging over many Monte Carlo runs.
@@
This stochastic sampling protocol is different conceptually from the collisional model approach. The microscopic modeling of the interaction and the partial tracing steps in the collisional models is replaced by proposing and then accepting or rejecting the jumps between the eigenstates. In other words, the system and the bath interaction followed by the partial tracing of the bath leads to the mixedness in the system density matrix in the collisional model picture. On other hand, in the Metropolis scheme the thermal density matrix is constructed from averaging over many runs. This has computational advantage, as the numerical steps involves dealing with Hilbert space of the system alone, while the same is not true in the collisional models. In the later approach, the combined system and the bath evolves under the interaction Hamiltonian, as mentioned in Eq. 1, thereby involving the computation over the combined Hilbert space.
@@
       Data: Initial eigenstate $\psi_{\rm ini}$
       for $i\leftarrow 1$ to $num\_runs$ do
             Initialize system into some randomly chosen eigenstate $\psi_{\rm ini}$;
@@
                   Propose a jump to a new eigenstate $\psi^{\prime}$;
@@
                         Accept the proposed eigenstate jump: $\psi\leftarrow\psi^{\prime}$;
@@
                         Reject the proposed eigenstate jump: $\psi$ remains unchanged;
                  Record the current state $\psi$.
      Calculate the thermal averages or occupation probabilities of the eigenstates by averaging over Monte Carlo runs.
@@
As an example, the thermalization of the XXZ model using the Metropolis technique is illustrated in Fig. 5. The parameters in the Hamiltonian are: $J=1,h=1,\Delta=1$ and $\beta=2$. As shown, the probabilities of the evolving state quickly approach the expected thermal probability values based on the thermal density matrix at the chosen inverse temperature.

### v2

@@ Abstract
Collisional models, or repeated interaction schemes, are a category of microscopic open quantum system models that have seen growing use in studying quantum thermalization, in which the bath is modeled as a large ensemble of identical ancillas that sequentially interact with the system. We demonstrate an equivalence between the system dynamics generated by the collisional model framework and the Metropolis algorithm, subject to two primary conditions. Namely, that each collisional model bath ancilla is prepared in a thermal state with a discrete spectrum that matches the energy eigenstate transitions of the system, and that the ratio of the ancilla partition function to the number of system eigenstates remains small. If these conditions are satisfied, the system dynamics generated by both methods are identical for arbitrary initial states and in both the steady state and transient regimes. This allows the typically purely phenomenological Metropolis scheme to be used as a tool to study exact pre-thermalization dynamics without the need to explicitly model the system-bath interaction.
@@ I Introduction
The mechanisms and conditions under which a many-body quantum system will thermalize is a question of significant interest that bridges the fields of quantum thermodynamics, condensed matter physics, atomic, molecular, and optical physics, and quantum information. In the context of open quantum systems it is well established that a quantum system coupled to a heat bath environment evolving under the Markovian Lindblad master equation will equilibrate to a thermal state at the temperature of the bath [1]. While the Lindblad equation is often tractable, it relies on strong assumptions about the dynamics of both the system and environment as well as the system-environment coupling, namely the Born-Markov and rotating wave approximations [1].
In recent years another approach to modeling open quantum systems has seen growing use, especially in the field of quantum thermodynamics, known as collisional models, or repeated interaction schemes [2]. In the collisional model approach the environment is assumed to consist of a collection of many identical subsystems, referred to as ancillae. The interaction between the system and environment occurs as a series of discrete unitary interactions (“collisions”) between the system and one ancilla of the environment.
By microscopically modeling the system-environment interaction, collisional models have proven particularly useful in studying non-Markovian dynamics [3, 4, 5, 6, 7, 8, 9], non-equilibrium dynamics [10, 11, 12, 13, 14, 15], quantum thermometry [16, 17], quantum synchronization [18, 19], strongly-correlated models [10, 20], quantum batteries and thermal machines [21, 22, 23], quantum optics [24, 25], and modeling noise in quantum devices [7]. Collisional models have also served as a successful framework for studying concepts at the intersection of information theory and thermodynamics such as information scrambling [26, 27], Landauer’s principle [28, 29], and quantum Darwinism [30, 31]. Notably, the typical Lindblad master equation can be derived from the collisional model framework under the assumptions of non-interacting, uncorrelated ancillae [32, 24, 33, 2].
@@
A notable drawback of the collisional model approach is the need to operate in the joint system-bath Hilbert space, which can become computationally unwieldy, especially for large-dimensional bath ancillae. Nevertheless, collisional models are conceptually important as they provide a microscopic framework that can operate outside of common assumptions such as weak system-bath interactions.
Distinct from the deterministic evolution in collisional models, Monte Carlo methods are another class of methods that rely upon stochastic sampling of states in the Hilbert space such that the dominant contributions to the ground or the thermal states are captured [40, 41]. These are iterative, Markovian algorithms wherein each iteration, a random change to the state is considered and either accepted or rejected based on a relative probability criterion. By stochastic averaging over different iterations, the underlying probability distribution can be built. The process must satisfy the principles of ergodicity and detailed balance, meaning that every state can be connected to any other state through a finite number of moves, to ensure a complete sampling of states. Such Monte Carlo techniques are utilized in understanding the thermodynamic properties of lattice systems [42, 43] and can also be interpreted as a dynamic process related to the Glauber Kinetic Ising models [44, 45].
Here we perform a detailed comparison between a typical collisional model approach of periodic system-bath interactions and a modified Monte Carlo Metropolis algorithm. We show that a Metropolis Monte Carlo scheme can replicate collisional model dynamics in both the transient and steady state regimes when the Metropolis time step is chosen to be the same as the time between collisions in the collisional model. Our equivalence holds under the assumption of thermalizing dynamics generated by interaction with a bath whose spectrum couples to each of the system’s energy eigenstate transitions. It is also subject to the condition that the ratio of the bath ancillae partition functions to the number of system energy eigenstate transitions is small, which itself depends on an interplay between temperature, model energy spectrum, and system size.
Notably, by modifying the rejection step procedure to implement a decoherence process the Metropolis algorithm, like the collisional model, can generate dynamics from arbitrary initial states, including coherent states. This equivalence shows that the Metropolis algorithm can be used not only as a phenomenological model for reaching thermalization, but can, under some circumstances, fully capture the transient dynamics of thermalization. Furthermore, as the Metropolis algorithm can be carried out without directly modeling the bath, this equivalence allows for thermalizing collisional model dynamics to be generated while only working in the Hilbert space of the system. While both methods still require diagonalizing of the system Hamiltonian to generate the transitions between energy eigenstates, the Metropolis algorithm posses a computational memory advantage over the collisional model since it only needs to store matrices that scale with the system Hilbert space, rather than the larger joint system plus bath Hilbert space.
@@ II Collisional model thermalization
This process, free evolution followed by interaction with a fresh environment ancilla, is then repeated for $n$ time steps. Note that the time interval between collisions $t_{S}=t_{n+1}-t_{n}$ depends on the physical properties of the system and environment, such as the system’s scattering cross section and the density of particles (ancillae) in the environment. In Fig. 1 we provide a conceptual illustration of the collisional model approach.
@@ II.1 Collisional model thermalization of the XXZ model
$H_{XXZ}=-J\sum_{q=1}^{N-1}\left(\sigma^{x}_{q}\sigma^{x}_{q+1}+\sigma^{y}_{q}\sigma^{y}_{q+1}+\Delta\sigma^{z}_{q}\sigma^{z}_{q+1}\right)+\frac{h}{2}\sum_{q=1}^{N}\sigma^{z}_{q}$ (4)
@@
$\begin{split}&E_{1}=-h-J\Delta,\quad\,E_{2}=J(\Delta-2),\\ &E_{3}=h-J\Delta,\qquad E_{4}=J(\Delta+2).\end{split}$ (5)
@@
We note that this is only one possible approach to constructing a bath that fulfills the energy matching condition. For example, another approach using low-dimensional ancillae with time-dependent energy gaps that are rapidly swept across the full system energy spectrum has also been demonstrated to lead to thermalization [46].
@@
Combining Eqs. (8), (9), and (6) in Eq. (1), we numerically simulate a 20 time step collisional model for the two-site XXZ chain. We assume the initial system state to be a pure state composed of an equal superposition of all energy eigenstates $\left|\psi_{0}\right\rangle=2^{-N/2}\sum_{i}\left|i\right\rangle$. We note there is nothing special about this choice of initial state, and thermalization will occur from an arbitrary initial state. In Fig. 3 we plot the occupation probabilities of each energy eigenstate as a function of the time step, $n$. We see that the occupation probabilities rapidly approach the thermal state values, indicating thermalization.
@@ III Thermalization under the Monte Carlo Metropolis algorithm
Next, we construct a Monte Carlo algorithm with Metropolis [47] updating to demonstrate thermalization in the same context as the collisional model. The Metropolis updating technique [47] was introduced as an algorithm to significantly improve Monte Carlo convergence. The goal of the Metropolis scheme is to generate a sequence of states such that the distribution of these states closely resembles the desired distribution. The key feature of this sampling procedure is using the Metropolis filter function when deciding to accept or reject a proposed move. There are two steps involved in this method. The first one is to propose a move from the present state $x$ to $x^{\prime}$, which is based on the conditional proposal probability $G(x^{\prime}|x)$. The next step involves the acceptance probability $A(x^{\prime}|x)$ which then determines the acceptance of the proposed move. The detailed balance condition requires, $A(x^{\prime}|x)G(x^{\prime}|x)P(x)=A(x|x^{\prime})G(x|x^{\prime})P(x^{\prime})$, where $P(x)$ is the underlying probability distribution to be sampled. Note that the normalization factor in $P(x)$ gets canceled and thus the stochastic averaging can be done without explicitly computing the normalization factor, one key advantage of the Metropolis method. In the context of thermalization, $P(x)$ is the Gibbs distribution, with the partition function as the normalization factor. Thus the accept and reject criteria are based on the energy difference between the proposed and the current state. For classical systems, for instance, spin systems, the updates are often local, involving the flipping of the spin at the chosen site. However, in general, the update scheme can also be non-local [48]. Since the spin basis states are the eigenstates of the classical spin Hamiltonians, these “classical” updates are sufficient to produce thermalization.
@@
In the context of quantum systems, the eigenvectors are, in general, non-trivial superpositions of the spin basis states. In this case, the classical spin updates are not sufficient to achieve thermalization. To do so, the update scheme must ensure the algorithm explores all the system eigenstates. This can be done using eigenstate jump operators, whose action produces a jump from the present eigenstate to any other eigenstate. The simplest choice for the proposal probability $G(x^{\prime}|x)$ is a uniform distribution, meaning that all possible “jumps” from the present eigenstate to other eigenstates will be proposed with equal probability.
@@
For studying thermalization, we consider an update scheme based on jumps between different eigenstates. The steps of the algorithm are outlined in Algorithm 1. The first step involves computing the eigenvectors $\{\psi_{i}\}$ and eigenvalues $\{E_{i}\}$ of the XXZ Hamiltonian given in Eq. 4 for a system of size $N$. The set of eigenvectors and eigenvalues can then be used for the Metropolis updating scheme. We consider an arbitrary initial state, $\rho_{0}$, which can be expressed in the energy eigenbasis as $\rho_{0}=\sum_{k,l}a_{k,l}\left|\psi_{k}\right\rangle\left\langle\psi_{l}\right|$. We then propose a transition $A_{i,j}=\left|j\right\rangle\left\langle i\right|$, with $j\neq i$, where $\left|i\right\rangle$ is an eigenstate with non-zero population in $\rho_{0}$ and $\left|j\right\rangle$ is chosen randomly. We then use the Metropolis condition based on the energy difference $E_{j}-E_{i}$ to decide whether the jump is accepted. If the jump is not accepted, we update the initial state to $\frac{1}{2}(\left|i\right\rangle\left\langle i\right|\rho_{0}+\rho_{0}\left|i\right\rangle\left\langle i\right|)$. This process serves to kill off the populations and coherences in $\rho_{0}$ other than those associated with $\left|i\right\rangle$, and acts as a discrete Monte Carlo implementation of the Lindblad decoherence term [1]. This step is then repeated many times, and the results are averaged over to construct the density matrix for the next time step. Repeating the entire sequence of proposal, accept/reject, and then averaging, $n$ times results in the construction of a time series that gives the time evolution of the density matrix. Note that across Monte Carlo runs for each time step coherences are only preserved by sequences of rejections and are reduced by half each time. Consequently, the coherences of the averaged time-dependent density matrix decrease exponentially, as expected in the decoherence process.
Our approach is related to the circuit-based quantum Metropolis scheme proposed in Ref. [49]. In this scheme a unitary operator is applied that rotates the initial eigenstate into a superposition state. If a transition is accepted, a measurement selects one element of this superposition state as the next eigenstate in the time series. Our Metropolis scheme takes a different approach by randomly choosing one out of all the potential eigenstate transitions, with equal probability given to each transition. Our algorithm also differs by directly incorporating decoherence during the rejection step. This allows states with non-zero coherences to be taken as an input to the algorithm, and (partially) preserves some coherences in the output time series. Other quantum implementations of Monte Carlo algorithms have also been proposed, such as the Monte-Carlo wave function method [50, 51, 52] and the quantum trajectories method [53] used in quantum optics.
@@
This stochastic sampling protocol in the Metropolis algorithm is different conceptually from the collisional model approach. The microscopic modeling of the interaction and the partial tracing steps in the collisional models is replaced by proposing and then accepting or rejecting the jumps between the eigenstates. In other words, the system and the bath interaction followed by the partial tracing of the bath leads to the mixedness in the system density matrix in the collisional model picture. On other hand, in the Metropolis scheme the thermal density matrix is constructed from averaging over many runs. This has computational advantage, as the numerical steps involves dealing with Hilbert space of the system alone, while the same is not true in the collisional models. In the later approach, the combined system and the bath evolves under the interaction Hamiltonian, as mentioned in Eq. 1, thereby involving the computation over the combined Hilbert space.
@@
       Data: Initial state $\rho_{0}$
       Initialize system into initial state $\rho_{0}$;
@@
             for $m\leftarrow 1$ to $num\_runs$ do
                   Propose jump $\left|\psi^{\prime}\right\rangle\left\langle\psi\right|$ where $\left|\psi\right\rangle$ is an eigenstate with non-zero population in $\rho$ and $\psi^{\prime}\neq\psi$;
@@
                         Accept the proposed jump: apply jump operator $\left|\psi^{\prime}\right\rangle\left\langle\psi\right|\rho\left|\psi\right\rangle\left\langle\psi^{\prime}\right|$;
@@
                         Reject the proposed jump: apply decoherence map $\frac{1}{2}(\left|\psi\right\rangle\left\langle\psi\right|\rho+\rho\left|\psi\right\rangle\left\langle\psi\right|)$;
            Average over Monte Carlo runs to generate current state $\rho^{\prime}$.
@@
As an example, the thermalization of the XXZ model using the Metropolis technique is illustrated in Fig. 5. The parameters in the Hamiltonian are: $J=1,h=1,\Delta=1$, and the inverse temperature is $\beta=2$. As in the case of the collisional model, we have taken the initial state to be a pure state composed of an even superposition of all the energy eigenstates. We can see the probabilities of the evolving state quickly approach the expected thermal probability values based on the thermal density matrix at the chosen inverse temperature.